Although noble gases normally do not react, xenon combines with elemental fluorine at elevated temperatures.

Xe + 2 F2 → XeF4

What is the theoretical mass of xenon tetrafluoride that should form when 130 g of xenon is reacted with 100 g of F2?




What is the percent yield if only 135 g of XeF4 is actually isolated?
1
Expert's answer
2020-12-16T05:54:57-0500

M (Xe) = 131.29 g/mol

M (XeF4) = 207.28 g/mol

n = m/M

n (Xe) = 130 / 131.29 = 1 mol

n (XeF4) = n (Xe) = 1 mol

m (XeF4) = n x M = 1 x 207.28 = 207 g

% (XeF4) = 135 / 207 x 100 = 65%